OpenStudy (anonymous):

A cube has side length 13 kilometers. What is its surface area?

OpenStudy (anonymous):

l*l where l is the side length. so the ans is 169km^2

OpenStudy (kainui):

First off, how many sides does a cube have? Try drawing one or looking at dice and counting the sides, or the sides of a book even. So once you know how many sides a cube has, you know that every side is made up of perfect squares, and so you can just find the area of one surface and multiply that by the total number of sides, there you go!

OpenStudy (anonymous):

312 km^2 1,014 km^2 2,197 km^2 156 km^2

OpenStudy (anonymous):

ssorry 6*l*l=136*6=ans

OpenStudy (anonymous):

Surface area = 6s^2

OpenStudy (anonymous):

169*6

OpenStudy (anonymous):

1014